Question 1: Triple Star Companys standard and actual costs per unit are provided below for the most recent period. During th i need answer of 4,5,6,7
Answer #1

Direct Labour Efficiencey Variance = Standard Rate X ( Standard time for actual output - Actual time worked)

= $ 7 X ( 3 - 3.2)

= $ 7 X (-0.2)

= - 1.4 (A-adverse)

Managerial Suggestion = Adverse for direct labour efficiency variance in above calculatin and for labour efficiency need of improvement. It will avoid this adverse situation in organisation.

Variable Overhead Rate Variance = ( Actual hrs. worked X Actual variable overhead rate per hrs.) - ( Actual hrs. worked X Standard Variable overhead rate per hrs.)

= ( 2.2 X $ 3.10) - ( 2.2 X $ 3)

= $ 6.82 - $ 6.6

= 0.22 (F-Favourable)

Managerial Suggestion = In abover calculation Variable Overhead Rate Variance is very positive for this variance as per estimated and it should be continue with this target.

Variable Overhead Efficiency Variance = (Actual labour hrs. - standard labour hrs.) X Standard overhead rate

= ( 2.2 - 2.0) X $ 3.0

= 0.2 X 3.0

= 0.6 (F-favourable)

Managerial Suggestion = From above calculation of Variable Overhead Efficiency Variance is favourable and its more efficience as per anticipated variance and it should be try to continue with this perfection.